Class A contains 3 boys and 5 girls, while class B contains 4 boys and 2 girls. If one person is selected from each class, the probability that both those chosen are girls is:
a. 5/24
b. 1/10
c. 23/24
d. 1/48
e. none of these

Answer by Fombitz(13828)

P(A,G)=5/8
P(B,G)=2/6=1/3
P(2 G)=(5/8)(1/3)=5/24
a) is the correct answer.